### 26 Feb 2024 v1

@@ Abstract
In the abstract Tile Assembly Model, self-assembling systems consisting of tiles of different colors can form structures on which colored patterns are “painted.” We explore the complexity, in terms of the unique tile types required, of assembling various patterns, proving several upper and lower bounds.
Keywords:
Self-assembly abstract Tile Assembly Model pattern assembly
@@ 1 Introduction
Systems in category (2), so-called fully-addressed or hard-coded have the benefit of being able to uniquely define each “pixel” of the structure and therefore to “paint” arbitrary pictures, which we’ll refer to as patterns, on their surfaces[15, 17, 11]. Although it is easy to see that any finite structure or pattern can self-assemble from a hard-coded system, it has previously been shown that there exist infinite structures and patterns that cannot self-assemble [8, 9]. The benefits of the algorithmic self-assembly[18, 5, 12] of category (3) include precise formation of shapes using exponentially fewer types of components[13, 14], thus reducing the cost, reducing the effort to fabricate and implement, and increasing the speed of growth of these systems[3]. However, the drastic reduction in the number of component types means a corresponding increase in their reuse. This results in copies of the same component type appearing in many locations throughout the resulting target structure, thus removing the ability to uniquely address each pixel when forming patterns. This generally results in a reduction in the number of patterns that are producible.
@@
In this paper we study the trade-off between the numbers of unique components, or tile types, needed to self-assemble designed patterns and the complexities of the patterns that can self-assemble. Our first results present constructions for making tile sets that self-assemble a series of relatively simple patterns to demonstrate how efficiently they can be built algorithmically. These consist of patterns of white and black pixels on the surfaces of squares, and include (1) a pattern with a single black pixel, (2) a pattern with some number $k$ of black pixels, and (3) grids of alternating black and white stripes. All of these are shown to have exponential reductions in tile type requirements, a.k.a. tile complexity, over fully-addressed systems.
@@
Our next pair of results combine to show a tight bound on the tile complexity of self-assembling arbitrary patterns of two colors on the surfaces of $n\times n$ squares for almost all $n$. Using an information theoretic argument we prove a lower bound, namely that for almost all $n$, such patterns have tile complexity $\Omega\left(\frac{n^{2}}{\log{n}}\right)$. We then provide a construction that, when given an arbitrary $n\times n$ pattern of black and white pixels, generates a tile set of $O\left(\frac{n^{2}}{\log{n}}\right)$ tile types that self-assemble an $n\times n$ square with black and white tiles that form that pattern. Although this is not a significant improvement over the $n^{2}$ tile types required to naively implement a fully-addressed set of tile types, the lower bound proves that for almost all patterns this is the best possible tile complexity.
@@
Although the prior result showed that any $n\times n$ pattern can self-assemble using $O\left(\frac{n^{2}}{\log{n}}\right)$ tile types, and in fact it is well known that any finite pattern or structure can self-assemble using a fully-addressed construction, our next result shows that, if given two planes in which tiles can self-assemble (one on top of the other) then it is possible for some patterns to self-assemble using asymptotically fewer tile types than when systems are restricted to a single plane. The proof of this result uses a novel application of diagonalization to tile-based self-assembly. Namely, one system simulates every system of a given tile complexity class for a bounded number of steps, sequentially, in order to algorithmically generate a pattern that is guaranteed not to be made by any of those systems, and then prints the pattern of that system on a second plane above the assembly that performed the simulations. We also show how to extend these square patterns infinitely to cover the plane, while maintaining the same tile complexity argument.
@@
Overall, our results demonstrate boundaries on tile complexities of algorithmic self-assembling systems when forming patterns, and help to demonstrate their benefits over fully-addressed systems. To make some of our results easier to understand, we have created a set of programs and tile sets that can be used to view examples. These can be found online: Pattern Self-Assembly Software [4].
@@ 2 Preliminary Definitions and Models
In this section we provide definitions for the terminology and model used throughout the paper.
@@ 2.1 The abstract Tile-Assembly Model
A tile-assembly system (TAS) is a triple $\mathcal{T}=(T,\sigma,\tau)$, where $T$ is a tile set, $\sigma$ is a finite $\tau$-stable assembly called the seed assembly, and $\tau\in\mathbb{Z}^{+}$ is called the binding threshold (a.k.a. temperature). Given a TAS $\mathcal{T}=(T,\sigma,\tau)$ and two $\tau$-stable assemblies $\alpha$ and $\beta$, we say that $\alpha$ $\mathcal{T}$-produces $\beta$ in one step (written $\alpha\to^{\mathcal{T}}_{1}\beta$) if $\alpha\sqsubseteq\beta$ and $|S_{\beta}\setminus S_{\alpha}|=1$. That is, $\alpha\to^{\mathcal{T}}_{1}\beta$ if $\beta$ differs from $\alpha$ by the addition of a single tile. The $\mathcal{T}$-frontier is the set $\partial^{\mathcal{T}}\alpha=\bigcup_{\alpha\to^{\mathcal{T}}_{1}\beta}S_{\beta}\setminus S_{\alpha}$ of locations in which a tile could $\tau$-stably attach to $\alpha$. When $\mathcal{T}$ is clear from context we simply refer to these as the frontier locations.
@@ 2.2 Patterns
Let $C$ be a set of colors and let $P\subseteq(\mathbb{Z}^{d}\times C)$. We say that $P$ is a $d$-dimensional pattern, i.e., a set of locations and corresponding colors. Let ${\rm dom}\;P$ to be the set of locations that are assigned a color. A pattern is k-colored when the number of unique colors is uses is $k$. Let $\texttt{Color}(P,\vec{l})$ be a function that takes a pattern and a location $\vec{l}$ and returns the color of the pattern at that location. (and is undefined if $\vec{l}\not\in{\rm dom}\;(P)$).
Given a TAS $\mathcal{T}=(T,\sigma,\tau)$, we allow each tile type to be assigned exactly one color from some set of colors $C$. Let $C_{P}\subseteq C$ be a subset of those colors, and $T_{C_{P}}\subseteq T$ be the subset of tiles of $T$ whose colors are in $C_{P}$. Given an assembly $\alpha\in\mathcal{A}[\mathcal{\mathcal{T}}]$, we use ${\rm dom}\;(\alpha)$ to denote the set of all locations with tiles in $\alpha$ and ${\rm dom}\;_{C_{p}}(\alpha)$ to denote the set of all locations of tiles in $\alpha$ with colors in $C_{P}$. Given a location $\vec{l}\in\mathbb{Z}^{d}$, let $\texttt{Color}(\alpha,\vec{l})$ also define a function that takes as input an assembly and a location and returns the color of the tile at that location (and is undefined if $\vec{l}\not\in{\rm dom}\;(\alpha)$). We say that $\mathcal{T}$ weakly self-assembles pattern $P$ iff for all $\alpha\in\mathcal{A}_{\Box}[\mathcal{T}]$, ${\rm dom}\;_{C_{P}}(\alpha)=P$ and $\forall(\vec{l},c)\in P,c=\texttt{Color}(\alpha,\vec{l})$. We say that $\mathcal{T}$ strictly self-assembles pattern $P$ iff $T_{C_{P}}=T$, i.e. all tiles of $T$ are colored from $C_{P}$, and $\mathcal{T}$ weakly self-assembles $P$ (i.e. all locations receiving tiles are within $P$).
Let $SQPATS_{n}$ be the set of all patterns that are $n\times n$ squares.
A pattern class is an infinite set of patterns parameterized by some values $X$. A pattern class can be represented as a function $PC:X\rightarrow P$ that maps parameters $X$ to some pattern $P$. A construction for a pattern class $PC$ is a function $C_{PC}:X\rightarrow T$ that takes the same parameters $X$ and outputs a TAS $T$ such that $T$ weakly self-assembles $PC(X)=P$.
@@ 3 Simple Patterns
In this section we define many different simple pattern classes and present constructions that can build those pattern classes.
@@
We define a pattern class that given an $n$, $i$, and $j$ creates a 2-colored $n\times n$ square white pattern with a single black tile located at $(i,j)$
$SINGLE\_PIXEL=(n,i,j)\rightarrow p$ where
  • $n,i,j\in\mathbb{N}$
  • $p\in\text{SQPATS}_{n}$
  • $n\geq i,j$
  • $\forall\vec{l}\in\{dom(p)-(i,j)\}\rightarrow\texttt{C}olor(p,\vec{l})=\text{ White }$
  • $\texttt{C}olor(p,(i,j))=\text{ Black }$
@@
For all $n,i,j\in\mathbb{N}$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$ and $\mathcal{T}$ strictly self-assembles $SINGLE\_PIXEL(n,i,j)$.
@@
We present a construction that given an n, i, j creates a TAS $T$ with a tile complexity of $O(\log n)$ that strictly self-assembles $P=SINGLE\_PIXEL(n,i,j)$. Figure 0(a) shows a high-level depiction of how we build the assembly.
The assembly starts by growing a hard-coded rectangle with an empty interior called a "counter box". Each edge of the box has glues for a counter to attach. The side lengths are $s=\log n$ this is the maximum length needed to count the entire square. Thus the counter box uses $O(\log n)$ tiles types.
Each side of the counter box has a number encoded in the glues. These numbers are pre-computed to grow a cross like structure that outlines the edges of the pattern. To complete the square, filler tiles will fill in between the counters and inside the counter box. There is a constant number of filler tiles.
The seed tile will be the single black tile. The counter box will grow from it. If the black tile is within $\log n$ of the side of the square, then the counter box will grow away from the edge. This means the counter box will always have room inside of the assembly.
We have shown that the tile complexity of this assembly is $O(n)$ and that it strictly self-assembles $P$. Thus the theorem is proved.
@@
We define a pattern class that given an $n$ and a list of locations $l$ with $\log n$ distance between each location creates a 2-colored $n\times n$ pattern with black tiles at each location in $l$ and white tiles everywhere else.
$MULTI-\_PIXEL=(n,l)\rightarrow p$ where
  • $n\in\mathbb{N}$
  • $l\in POWERSET({\rm dom}\;p)$
  • $p\in SQPATS_{n}$
  • $\forall\vec{v}\in dom(p)\rightarrow\texttt{C}olor(p,\vec{v})=$ Black if $\vec{v}\in l$ else White
  • $\forall(x,y)\in l\rightarrow\forall(x^{\prime},y^{\prime})\in l\rightarrow|x-x^{\prime}|\geq\log n\land|y-y^{\prime}|\geq\log n$
Figure 2: (a) An example of a multi-pixel pattern with 3 black pixels. (b) A tree of counters is constructed to go to each pixel and the edges of the box. The counter box is colored red. The counters are colored blue. The white locations are filled by generic filler tiles.
@@
For all $n\in\mathbb{N}$ and sets of locations $l$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$ and $\mathcal{T}$ strictly self-assembles $MULTI\_PIXEL(n,l)$.
Proof
We present a construction that given an n, l, will provide a TAS $T$ with a tile complexity of $O(\log n)$ that strictly self-assembles $MULTI\_PIXEL(n,l)$. Figure 1(a) shows a high-level depiction of how we build the assembly.
We first build a path that visits each pixel in $l$. The path can only grow horizontally and vertically. The path can not intersect itself but it can branch anywhere. This can be done for any set of points. Every point where the path changes directions or there is a black pixel is called a node. The path must also touch each edge of the square.
We set the seed to be the tile at the beginning of the path. Each segment of the path has a unique counter that counts the length of the segment and builds the next counter. Every node on the path builds a counter box that counts to the next nodes in the path. The nodes that correspond to black pixels will paint the corresponding corner of them self black to adhere to the pattern. Since the points are all $\log n$ distance away, the counter boxes will never overlap.
Filler tiles fill inside the counter boxes and in-between the segments of the path. Since the path touches all the edges of the square, the entire square is filled, forming the entire pattern
Each counter box uses $O(\log n)$ tiles types. There are a constant $|l|$ number of counter boxes. There is a constant number of filter tiles. Thus the tile complexity of $T$ is $O(\log n)$ and the theorem is proved.
@@
We define a pattern class that given an $n$, $i$, and $j$ creates a 2-colored $n\times n$ square that has black stripes every $i$ rows and $j$ columns
$STRIPES=(n,i,j)\rightarrow p$ where
  • $n\in\mathbb{N}$
  • $i,j\in\mathbb{N}$
  • $i,j<n$
  • $p\in SQPATS_{n}$
  • $\forall x\in[n]\rightarrow\forall y\in[n]\rightarrow\texttt{C}olor(p,(x,y))=\text{Black}\text{ if }x\mod i=0\text{ or }y\mod j=0\text{ else White}$
@@
For all $n,i,j\in\mathbb{N}$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$ and $\mathcal{T}$ strictly self-assembles $STRIPES(n,i,j)$.
Proof
We present a construction that given an n, i, and j, outputs a $\mathcal{T}$ that strictly self assembles $STRIPES(n,i,j)$. Figure 2(a) shows a high-level depiction of how we build the assembly.
We start by creating four counters. One of the counters counts upwards to $i$ and then place a stripe. Another counter grows alongside the first counter and only increments after each stripe. This counter counts the number of stripe in the assembly which equals $\frac{n}{i}$. The other two counters do the same for the other directions. The stripes will grow along side the counters and will go from one end of the square to the other. The counters are only in the first row and column they count out forming an L shape. The stripe tiles and the white filler tiles will fill the rest of the square with the pattern.
This construction needs 4 hard coded counters that each use $O(\log n)$ tile types. There is a constant number of filler and stripe tiles. The seed tile is a single tile that will grow the starting values for all the counter. Thus the theorem is proved.
Figure 3: (a) An example of a stripes pattern. (b) The blue tiles count to the next stripe while the red tiles count the number of stripes.

### 10 Mar 2024 v2

@@ Abstract
In the abstract Tile Assembly Model, self-assembling systems consisting of tiles of different colors can form structures on which colored patterns are “painted.” We explore the complexity, in terms of the numbers of unique tile types required, of assembling various patterns.We first demonstrate how to efficiently self-assemble a set of simple patterns, then show tight bounds on the tile type complexity of self-assembling 2-colored patterns on the surfaces of square assemblies. Finally, we demonstrate an exponential gap in tile type complexity of self-assembling an infinite series of patterns between systems restricted to one plane versus those allowed two planes.
@@ 1 Introduction
Systems in category (2), so-called fully-addressed or hard-coded have the benefit of being able to uniquely define each “pixel” of the structure and therefore to “paint” arbitrary pictures, which we’ll refer to as patterns, on their surfaces[21, 23, 17]. Although it is easy to see that in models of tile-based self-assembly such as the abstract Tile Assembly Model (aTAM)[22] any finite structure or pattern can self-assemble from a hard-coded system, it has previously been shown that there exist infinite structures and patterns that cannot self-assemble in the aTAM [12, 13]. The benefits of the algorithmic self-assembly[24, 8, 18] of category (3) include precise formation of shapes using exponentially fewer types of components[19, 20], thus reducing the cost, reducing the effort to fabricate and implement, and increasing the speed of growth of these systems[5]. However, the drastic reduction in the number of component types means a corresponding increase in their reuse. This results in copies of the same component type appearing in many locations throughout the resulting target structure, thus removing the ability to uniquely address each pixel when forming patterns. This generally results in a reduction in the number of patterns that are producible.
@@
In this paper we study the trade-off between the numbers of unique components, or tile types, needed to self-assemble designed patterns and the complexities of the patterns that can self-assemble. Past lines of work have dealt with the self-assembly of patterns in the aTAM from the perspective of the computational complexity of designing minimal tile sets to self-assemble given patterns (the so-called “PATS” problem)[11, 4, 15, 14] and others have shown the possibilities and impossibilities of assembling some classes of infinite patterns[12, 16]. In contrast, our first results present constructions for making tile sets that self-assemble a series of relatively simple patterns to demonstrate how efficiently they can be built algorithmically. These consist of patterns of white and black pixels on the surfaces of squares, and include (1) a pattern with a single black pixel, (2) a pattern with some number $k$ of black pixels, and (3) grids of alternating black and white stripes. All of these are shown to have exponential reductions in tile type requirements, a.k.a. tile complexity, over fully-addressed systems.
@@
Our next pair of results combine to show a tight bound on the tile complexity of self-assembling arbitrary patterns of two colors on the surfaces of $n\times n$ squares for almost all patterns. Using an information theoretic argument we prove a lower bound, namely that for almost all patterns on $n\times n$ squares, such patterns have tile complexity $\Omega\left(\frac{n^{2}}{\log{n}}\right)$. We then provide a construction that, when given an arbitrary $n\times n$ pattern of black and white pixels, generates a tile set of $O\left(\frac{n^{2}}{\log{n}}\right)$ tile types that self-assemble an $n\times n$ square with black and white tiles that form that pattern. Although this is not a significant improvement over the $n^{2}$ tile types required to naively implement a fully-addressed set of tile types, the lower bound proves that for almost all patterns this is the best possible tile complexity.
@@
Although the prior result showed that any $n\times n$ pattern can self-assemble using $O\left(\frac{n^{2}}{\log{n}}\right)$ tile types, our next result shows that, if given two planes in which tiles can self-assemble (one on top of the other) then it is possible for some patterns to self-assemble using exponentially fewer tile types than when systems are restricted to a single plane. (In fact, this result can be modified for arbitrary separation in tile complexity.) The proof of this result uses a novel application of diagonalization to tile-based self-assembly. Namely, one system simulates every system of a given tile complexity class for a bounded number of steps, sequentially and within a square of one plane, in order to algorithmically generate a pattern that is guaranteed not to be made by any of those systems, and then prints that generated pattern on the square of the second plane above the assembly that performed the simulations. We also show how to extend these square patterns infinitely to cover the plane, while maintaining the same tile complexity argument.
@@
Overall, our results demonstrate boundaries on tile complexities of algorithmic self-assembling systems when forming patterns, and help to demonstrate their benefits over fully-addressed systems. To make some of our results easier to understand, we have created a set of programs and tile sets that can be used to view examples. These can be found online: Pattern Self-Assembly Software [7]. Due to space limitations, proofs have been moved to a Technical Appendix in this version, and this full version can also be found online [6].
@@ 2 Preliminary Definitions and Models
In this section we define the terminology and model used throughout the paper.
@@ 2.1 The abstract Tile-Assembly Model
A tile-assembly system (TAS) is a triple $\mathcal{T}=(T,\sigma,\tau)$, where $T$ is a tile set, $\sigma$ is a finite $\tau$-stable assembly called the seed assembly, and $\tau\in\mathbb{Z}^{+}$ is called the binding threshold (a.k.a. temperature). If the seed $\sigma$ consists of a single tile, i.e. $|\sigma|=1$, we say $\mathcal{T}$ is singly-seeded. Given a TAS $\mathcal{T}=(T,\sigma,\tau)$ and two $\tau$-stable assemblies $\alpha$ and $\beta$, we say that $\alpha$ $\mathcal{T}$-produces $\beta$ in one step (written $\alpha\to^{\mathcal{T}}_{1}\beta$) if $\alpha\sqsubseteq\beta$ and $|S_{\beta}\setminus S_{\alpha}|=1$. That is, $\alpha\to^{\mathcal{T}}_{1}\beta$ if $\beta$ differs from $\alpha$ by the addition of a single tile. The $\mathcal{T}$-frontier is the set $\partial^{\mathcal{T}}\alpha=\bigcup_{\alpha\to^{\mathcal{T}}_{1}\beta}S_{\beta}\setminus S_{\alpha}$ of locations in which a tile could $\tau$-stably attach to $\alpha$. When $\mathcal{T}$ is clear from context we simply refer to these as the frontier locations.
@@ 2.2 Patterns
Let $C$ be a set of colors and let $P\subseteq(\mathbb{Z}^{d}\times C)$. We say that $P$ is a $d$-dimensional pattern, i.e., a set of locations and corresponding colors. Let ${\rm dom}\;P$ be the set of locations that are assigned a color. A pattern is k-colored when the number of unique colors used is $k$. Let $\texttt{Color}(P,\vec{l})$ be a function that takes a pattern and a location $\vec{l}$ and returns the color of the pattern at that location. (and is undefined if $\vec{l}\not\in{\rm dom}\;(P)$).
Given a TAS $\mathcal{T}=(T,\sigma,\tau)$, we allow each tile type to be assigned exactly one color from some set of colors $C$. Let $C_{P}\subseteq C$ be a subset of those colors, and $T_{C_{P}}\subseteq T$ be the subset of tiles of $T$ whose colors are in $C_{P}$. Given an assembly $\alpha\in\mathcal{A}[\mathcal{\mathcal{T}}]$, we use ${\rm dom}\;(\alpha)$ to denote the set of all locations with tiles in $\alpha$ and ${\rm dom}\;_{C_{p}}(\alpha)$ to denote the set of all locations of tiles in $\alpha$ with colors in $C_{P}$. Given a location $\vec{l}\in\mathbb{Z}^{d}$, let $\texttt{Color}(\alpha,\vec{l})$ define a function that takes as input an assembly and a location and returns the color of the tile at that location (and is undefined if $\vec{l}\not\in{\rm dom}\;(\alpha)$). We say $\mathcal{T}$ weakly self-assembles pattern $P$ iff for all $\alpha\in\mathcal{A}_{\Box}[\mathcal{T}]$, ${\rm dom}\;_{C_{P}}(\alpha)=P$ and $\forall(\vec{l},c)\in P,c=\texttt{Color}(\alpha,\vec{l})$. We say $\mathcal{T}$ strictly self-assembles pattern $P$ iff $T_{C_{P}}=T$, i.e. all tiles of $T$ are colored from $C_{P}$, and $\mathcal{T}$ weakly self-assembles $P$ (i.e. all locations receiving tiles are within $P$).
Let the set of all patterns be $\mathbb{P}$, the set of all $c$-colored patterns be $\mathbb{P}_{c}$, and $\texttt{SQPATS}_{c,n}\subset\mathbb{P}_{c}$ be the set of all $c$-colored patterns that are on the surfaces of $n\times n$ squares. Let $\texttt{SQPATS}_{c}=\bigcup_{n\in\mathbb{Z}^{+}}{\texttt{SQPATS}_{c,n}}$, i.e., the set of all $c$-colored patterns that are on the surfaces of squares. A pattern class is an infinite set of patterns parameterized by some set of values $X$, and can be represented as a function $PC:X\rightarrow\mathbb{P}$ that maps parameters $X$ to some pattern $P\in\mathbb{P}$. Let $\mathbb{T}$ be the set of all aTAM systems. A construction for pattern class $PC$ is a function $C_{PC}:\mathbb{P}\rightarrow\mathbb{T}$ that takes a pattern $P\in\mathbb{P}$ and outputs an aTAM system $\mathcal{T}\in\mathbb{T}$ such that $\mathcal{T}$ weakly self-assembles $P$.
@@ 3 Simple Patterns
In this section, we define several relatively simple pattern classes and present constructions that can build them efficiently.
First we define a pattern class whose patterns each consist of a 2-colored $n\times n$ square that is completely white except for a single black pixel.
@@
Given $n,i,j\in\mathbb{N}$, where $i,j<n$, define $\texttt{SinglePixel}(n,i,j)\rightarrow P$ such that (1) $P\in\texttt{SQPATS}_{2,n}$, (2) $\forall\vec{l}\in dom(P)-\{(i,j)\}$, $\texttt{C}olor(P,\vec{l})=\text{White}$, and (3) $\texttt{C}olor(P,(i,j))=\text{Black}$.
@@
For all $n,i,j\in\mathbb{N}$ such that $n\geq i,j$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$, and $\mathcal{T}$ weakly self-assembles $\texttt{SinglePixel}(n,i,j)$.
@@
We present a construction that, given an $n,i,j\in\mathbb{N}$ such that $i,j<n$, creates a TAS $\mathcal{T}$ with tile complexity of $O(\log n)$ that weakly self-assembles $P=\texttt{SinglePixel}(n,i,j)$. Figure 1 shows a high-level depiction of how we build the assembly.
The assembly starts by growing a hard-coded rectangle with an empty interior called a counter box. Each side of the box has glues for a counter to attach. The side lengths are $s=\lceil\log n\rceil$, as this is the maximum length needed to encode the bits required for a binary counter to count a full dimension of the entire square (which is the maximum that could be necessary). Thus, the counter box uses $O(\log n)$ tiles types.
Each side of the counter box has a number encoded in the outward-facing glues. These numbers are pre-computed so that binary counter tiles (i.e. sets of tiles that operate as standard binary counters) grow outward from them to form a cross-like structure that extends to each boundary of the $n\times n$ square. A constant-sized set of tile types (independent of $i$,$j$, and $n$) is used for each of the four counters. To complete the $n\times n$ square, a constant-sized set of filler tiles fill in between the counters and inside the counter box.
The seed tile is a single black tile, from which the counter box grows, while all other tile types are white. If the black tile is within $\log n$ of the side of the square, then the counter box grows away from the edge, meaning the counter box always has room to grow inside of the $n\times n$ square.
We have shown that the tile complexity of this assembly is $O(\log n)$ and that it weakly self-assembles $P$. Thus, the Theorem 3.1 is proved.
@@
Next, we define a pattern class whose patterns each consist of a 2-colored $n\times n$ square that is completely white except for a set of (separated and individual) black pixels.
@@
Given $n\in\mathbb{N}$ and a set of locations $L\subseteq[0,n-1]^{2}$ such that $\forall(x,y),(x^{\prime},y^{\prime})\in L$, $|x-x^{\prime}|\geq\lceil\log n\rceil\lor|y-y^{\prime}|\geq\lceil\log n\rceil$, define $\texttt{MultiPixel}(n,L)\rightarrow P$ such that (1) $P\in\texttt{SQPATS}_{2,n}$, and (2) $\forall\vec{v}\in dom(P),\texttt{C}olor(P,\vec{v})=$ Black if $\vec{v}\in L$ else White.
Figure 2: (a) An example of a multi-pixel pattern with three black pixels. (b) A tree of counters is constructed to grow to each pixel and the edges of the square. The counter boxes are colored red. The counters are colored blue. The white locations are filled by generic filler tiles.
@@
For all $n\in\mathbb{N}$ and sets of locations $L\subseteq[0,n-1]^{2}$ such that $\forall(x,y),(x^{\prime},y^{\prime})\in L$, $|x-x^{\prime}|\geq\lceil\log n\rceil\lor|y-y^{\prime}|\geq\lceil\log n\rceil$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(|L|\log n)$, and $\mathcal{T}$ weakly self-assembles $\texttt{MultiPixel}(n,L)$.
The proof of Theorem 3.2 uses a construction that is an extension of that used in the proof of Theorem 3.1, and essentially uses a series of counters to build a path connecting all pixels and filler tiles for the remaining portion of the square. A high-level depiction is shown in Figure 2, and full details can be found in Section 7.1.
The final (relatively) simple pattern class that we define contains patterns that each consist of a $2$-colored $n\times n$ square with a set of repeating black horizontal rows and a set of repeating black vertical columns.
@@
Given $n,i,j\in\mathbb{N}$, where $i,j<n$, define $\texttt{Stripes}(n,i,j)\rightarrow P$ such that (1) $P\in\texttt{SQPATS}_{2,n}$, and (2) $\forall x,y\in[0,n-1],\texttt{C}olor(P,(x,y))=\text{Black}\text{ if }x\mod i=0\text{ or }y\mod j=0,\text{ else White}$.
@@
For all $n,i,j\in\mathbb{N}$, where $i,j<n$, there exists an aTAM system $\mathcal{T}=(T,\sigma,2)$ such that $|\sigma|=1$, $|T|=O(\log(n))$ and $\mathcal{T}$ weakly self-assembles $\texttt{Stripes}(n,i,j)$.
The proof of Theorem 3.3 can be found in Section 7.2. It is done by construction, where the construction has counters that grow vertically to count to, and mark, the locations of horizontal strips, and counters that grow horizontally to count to, and mark, the locations of vertical stripes. Counters also keep track of the distance to the boundaries of the square to ensure growth stops at the correct locations. An overview can be seen in Figure 3.
Figure 3: (a) An example of a stripes pattern. (b) The blue tiles count to the next stripe, while the red tiles count the number of stripes. Green tiles represent the starting rows for the counters (with the seed tile being at the corner where the green row and column intersect). Dark grey tiles represent counter tiles that are colored black